OpenStudy (justonemathproblem):

Just one question that's got me stuck out of a 105 question test....

OpenStudy (justonemathproblem):

Charlie was carrying around the steps to solve Square root of 6x plus 4 = 8. On his way through the Math Factory, Charlie dropped them and they mixed with other steps. Put the correct steps in order to solve Square root of 6x plus 4 = 8. A. Divide by 2 from both sides B. Square root both sides C. Divide by 6 from both sides D. Subtract 4 from both sides E. Subtract 8 from both sides F. Square both sides

OpenStudy (justonemathproblem):

options: A- First A, then F, then E B- First B, then D, then C C- First F, then D, then C

OpenStudy (justonemathproblem):

@callielovesyhuu

jhonyy9 (jhonyy9):

how is exactly your exercise wrote sqrt(6x +4) =8 or sqrt6x +4 =8 ?

OpenStudy (mathstudent55):

Did you ever notice that above the equal sign there is a plus sign on the keyboard? Is this the problem? \(\sqrt{6x + 4} = 8\)

OpenStudy (mathstudent55):

If so, square both sides first. 6x + 4 = 64 Subtract 4 from both sides. 6x = 60 Divide both sides by 6. x = 10